OpenStudy (anonymous):

find the equation of the tangent line to sin(x+y)=xy at (x,y)=(0,0)

OpenStudy (cwrw238):

first differentiate implicitly

OpenStudy (cwrw238):

cos(x + y)(1 + dy/dx) = x dy/dx + y now find dy/dx in terms of x and y then plug in x = 0 and y = 0 to get the slope of the tangent line.

OpenStudy (anonymous):

so in term of X. the final result should be y=cos(x+y)-y/x-cos(x+y)

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

cos(x + y) + cos(x + y)dy/dx= x dy/dx + y cos(x + y)dy/dx - x dy/dx = y - cos(x+y) dy/dx = y - cos(x+y) ------------ cos(x + y) - x x = 0 and y = 0 dy/dx = 0 - cos 0 -------- = -1 cos 0 - 0 so the slope of the tangent line is -1 y - y1 = m(x - x1) x1 = y1 = 0 , m = -1 y = -x is the required tangent line